OpenStudy (maddieayeeee):

What is the volume of a coffee can that is 6.5 inches high and has a diameter of 5 inches? Round your answer to the nearest tenth.

563blackghost (563blackghost):

Use this formula. \(\huge\bf{V=\pi r ^{2} h}\) We must first identify the radius so we would divide the diameter by 2. \(\huge\bf{5 \div 2 =?}\)

OpenStudy (maddieayeeee):

2.5

563blackghost (563blackghost):

Nice now we input this into the equation. \(\huge\bf{V= 3.14(2.5)^{2}(6.5)}\) We first simplify the paranthesis.... \(\huge\bf{V= 3.14 \times 6.25 \times 6.5}\) What would it equal?

OpenStudy (maddieayeeee):

127.5625

OpenStudy (maddieayeeee):

@563blackghost

563blackghost (563blackghost):

Correct :)